5. Find the kinetic energy of an electron whose de Broglie wavelength is 12.0 nm.
Answer #1

\lambda =\frac{h}{\sqrt{2*m*(Kinetic\ Energy)}}

Kinetic\ Energy=\frac{h^{2}}{2m\lambda ^{2}}=\frac{(6.626*10^{-34})^{2}}{2*9.1*10^{-31}*(12*10^{-9})^{2}}= 1.675*10^{-21}\ J